6(x-3)=x+8
a. x=14/5
b. x=26/5
c. x=13/3
d. x=11/5
I worked this problem, but I am not for sure if I worked it right.
But, I chose Letter "B"
Please help

Answer by checkley77(12844)

6(x-3)=x+8
6x-18=x+8
6x-x=8+18
5x=26
x=26/5 ans. B).